OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve x^3+y^3=y+21 at the point ( 3 , -2 ) and use it to find the equation of the tangent line in the form y=mx+b.

myininaya (myininaya):

\[(x^3)'=?\]

OpenStudy (anonymous):

3x^2

myininaya (myininaya):

\[(y^3)'=?\]

OpenStudy (anonymous):

3y^2

myininaya (myininaya):

Now exactly

myininaya (myininaya):

Not*

OpenStudy (anonymous):

3y^2 y'

myininaya (myininaya):

\[3x^2+3y^2y'=y'+0\]

myininaya (myininaya):

Solve for y'

OpenStudy (anonymous):

(1-3x^2)/(3y^2)

myininaya (myininaya):

\[3yy'-y'=-3x^2\]

myininaya (myininaya):

See how I got my y' terms together

OpenStudy (anonymous):

yes, and then y' turns into -1 or you do put it in front like y'[3y]

myininaya (myininaya):

I did that so I can factor out y' and divide by whatever it is being multiplied by

myininaya (myininaya):

\[y'(3y^2-1)=-3x^2\]

myininaya (myininaya):

\[y'=\frac{-3x^2}{3y^2-1}\]

OpenStudy (anonymous):

and then i just plug in (3,-2) for x and y?

myininaya (myininaya):

yes 3 for x and -2 for y

OpenStudy (anonymous):

Thank you very much, the answer was y=-27/11x+59/11

myininaya (myininaya):

Great stuff! :) I'm glad you got it!

OpenStudy (anonymous):

another one that was giving me trouble was this one Use implicit differentiation to find the slope of the tangent line to the curve xy^3+xy= 14 at the point ( 7 , 1 ). m=

myininaya (myininaya):

\[(xy^3)'=? \] You need product rule and chain rule

myininaya (myininaya):

\[=(x)'y^3+x(y^3)'\]

myininaya (myininaya):

\[(xy^3)'=1 \cdot y^3+x \cdot 3y^2 y'\]

myininaya (myininaya):

You try (xy)'

OpenStudy (anonymous):

1*y+xy'

myininaya (myininaya):

ok great!

myininaya (myininaya):

so we have \[ y^3+x \cdot 3y^2 y'+y+xy'=0\]

OpenStudy (anonymous):

is it y'=(-y^3-y)/(3xy^2+x)

myininaya (myininaya):

yep :)

OpenStudy (anonymous):

m=-1/14 woooo thanks again :)

myininaya (myininaya):

all you! :)

OpenStudy (anonymous):

Here's another where it uses log and I don't know what to do. Use implicit differentiation to find the slope of the tangent line to the curve 3^x+\log_2(xy)=10 at the point ( 2 , 1 ) and use it to find the equation of the tangent line in the form y=mx+b.

OpenStudy (anonymous):

3^x+log_2(xy)=10

myininaya (myininaya):

\[(a^x)'=\ln(a) \cdot a^x\] a>0 ---- \[(log_2(xy))'=(\frac{\ln(xy)}{\ln(2)})'=\frac{1}{\ln(2)} \cdot (\ln(xy))'=\frac{1}{\ln(2)} \cdot \frac{(xy)'}{xy}\]

OpenStudy (anonymous):

where did the ln go in (ln(xy))′

myininaya (myininaya):

(ln(f))'=f'/f (ln(inside))'=derivative of inside/inside

OpenStudy (anonymous):

Is this right? 1/ln(2) * (1*y'+x*1)/(xy)

myininaya (myininaya):

(xy)'=y+xy'

OpenStudy (anonymous):

Ahh oops, I just saw that

myininaya (myininaya):

Everything else looks good :)

myininaya (myininaya):

So we have: \[\ln(3) \cdot 3^x+\frac{1}{\ln(2)} \cdot \frac{y+xy'}{xy}=0\] agree?

OpenStudy (anonymous):

why did you take the ln of 3^x?

myininaya (myininaya):

\[a>0,(a^x)'=\ln(a) \cdot a^x\] \[y=a^x\] Take ln( ) of both sides! \[\ln(y)=\ln(a^x)\] Use one of your properties to rewrite \[\ln(y)=x \ln(a)\] Now differentiate both sides! \[\frac{y'}{y}=\ln(a)\] Multiply y on both sides to isolate y' So we have \[y'= y \ln(a)\] \[ \text{ recall from the beginning that } y=a^x\] So we have \[y'=a^x \cdot \ln(a) \]

OpenStudy (anonymous):

Okay, I see it now ln(3)⋅3x+1ln(2)⋅y+xy′xy=0

OpenStudy (anonymous):

3^x*

myininaya (myininaya):

\[\ln(3) \cdot 3^x+\frac{1}{\ln(2)} \cdot \frac{y+xy'}{xy}=0 ?\]

OpenStudy (anonymous):

yes that, apparently copy paste doesn't work..it left out the division signs

myininaya (myininaya):

Ok so we need to isolate y' , right? I would start by putting terms without the factor y' on the opposite side of terms the include the factor y'

myininaya (myininaya):

that include the factor y'*

myininaya (myininaya):

Like so... \[\frac{1}{\ln(2)} \cdot \frac{y+xy'}{xy}=-\ln(3) \cdot 3^x\]

myininaya (myininaya):

Now multiply both sides by the left hand's bottom's junk!

OpenStudy (anonymous):

like this? y+xy'=(-ln(3)*3^x)/(ln(2)xy)

myininaya (myininaya):

You multiplied on one side but divided on the other by the junk I told you to multiply on both sides

OpenStudy (anonymous):

y+xy'=(-ln(3)*3^x)(ln(2)xy)

myininaya (myininaya):

Perfect! :)

OpenStudy (anonymous):

y'=((-ln(3)*3^x)(ln(2)xy)-y)/(x)

myininaya (myininaya):

Yep yep Great job kolo!

myininaya (myininaya):

Ok hey I have to go Do you think you got it from here? If not, please post a link in the chats? I'm sorry.

OpenStudy (anonymous):

Yes, I think I can take it from here, thanks so much for all the help!

myininaya (myininaya):

No problem. Sorry I have to go :( It was fun differentiating stuff with you!

OpenStudy (anonymous):

Haha indeed it was, thanks again :)